In Econoland, let the amount of consumption that is independent of income is equal to 80; all expenditures are in billion US dollars. Investment in the economy is autonomous at 600. From any dollar change in national income, households use up ¾ of it to spend on consumption. Government spending is fixed at 300, and the tax rate is known to be 5%.

The correct statement based on this case is

Group of answer choices

a. multiplier is 4

b. After accounting for the tax rate, the marginal propensity to consume is 0.7125

c. Marginal propensity to import is 0.75

d. Marginal propensity to save is 0.25

e. The consumption function is C = 600 + 0.75Y
Answer #1

From any dollar change in national income, households use up ¾ of it to spend on consumption.

It means, MPC = 3/4

MPC = 0.75

MPS + MPC = 1

MPS = 1 - MPC

MPS = 1-0.75

MPS = 0.25

tax rate = 5% = 0.05

MPC after accounting tax rate = MPC (1- 0.05)

MPC after accounting tax rate = 0.75 (0.95)

MPC after accounting tax rate = 0.7125

Multiplier = 1 / [1 - MPC (1-t)]

Multiplier = 1 / [1 - 0.75(1-0.05)]

Multiplier = 1 / [1 - 0.7125]

Multiplier = 3.478

Autonomous consumption = 80

Consumption function: C = 80 + 0.75Yd

Answer: Option (b) and Option (d)